OpenStudy (sparklestaraa):

Winnie poured 14 cups of water into a rectangular container measuring 13 inches by 7 inches by 6 centimeters. [1 cup = 14.44 cubic inches; 1 inch = 2.54 cm] Part A: What is the maximum volume of water that the rectangular container can hold? Show your work. Part B: How many cubic inches of water was poured into the container? Part C: What was the height of the water in the container? Show your work.

OpenStudy (sparklestaraa):

@dtan5457 This i need help on just a little to understand

OpenStudy (dtan5457):

For starters part A: just use the formula for a regular rectangular prism which is length times width times height

OpenStudy (dtan5457):

convert the 6 cm to inches first

OpenStudy (dtan5457):

6cm=2.36220472441 which you might want to round

OpenStudy (dtan5457):

multiply 2.36220472441,13,7

OpenStudy (dtan5457):

for volume in inches

OpenStudy (dtan5457):

PART B: 1 cup=14.44 cubic inches.. so if they poured 14 of them... 14.44 times 14=202.16 cubic inches

OpenStudy (dtan5457):

as for part c: 202.16 = 13*7*h 202.16 from the water cubic inches, now knowing your other dimensions, figure out the height with that equation

OpenStudy (dtan5457):

@Sparklestaraa

OpenStudy (sparklestaraa):

:D I'm reading it now @dtan5457

OpenStudy (sparklestaraa):

10.64 is the height?? @dtan5457

OpenStudy (dtan5457):

202.16 = 13*7*h 13 times 7=91 divide 202.16 by 91

OpenStudy (dtan5457):

for your height ^^

OpenStudy (sparklestaraa):

2.22?

OpenStudy (sparklestaraa):

@dtan5457

OpenStudy (dtan5457):

yes

OpenStudy (sparklestaraa):

:D thanks so much i'd give a million medals if i could :D can u help with a few more??

OpenStudy (dtan5457):

sure. just tag me onto a new question.